Harrangue Company's standard variable overhead rate is $6 per direct labor hour, and each unit requires 2 standard direct labor hours. During March, Harry recorded 6,000 actual direct labor hours, $37,000 actual variable overhead costs, and 2,900 units of product manufactured.

What is the variable overhead efficiency variance for March for Harrangue?

a.$2,200 (U)

b.$1,200 (U)

c.$600 (U)

d.$2,200 (F)

Synergy Manufacturing Company has a normal monthly activity of 7,500 units.

Standard factory overhead rates are based on a normal monthly volume of one standard direct hour per unit.

Standard factory overhead rates per direct labor hour are:

Fixed $ 5.00
Variable 12.00 $17.00
Units actually produced in current month 7,000 units
Actual factory overhead costs incurred (includes $50,000 fixed) $140,000


What is the variable overhead spending variance for Synergy?

a.$6,000 unfavorable

b.$21,000 unfavorable

c.$0

d.$6,000 favorable

Malkovich Company uses a standard costing system. The following information pertains to direct materials for the month of July:

Standard price per lb. $18.00
Actual purchase price per lb. $16.50
Quantity purchased 3,100 lbs.
Quantity used 2,950 lbs.
Standard quantity allowed for actual output 3,000 lbs.
Actual output 1,000 units


Malkovich Company reports its material price variances at the time of purchase.

What is the material usage variance for Malkovich Company?

a.$900 (F)

b.$1,950 (F)

c.$900 (U)

d.$2,850 (F)

Bender Corporation produced 100 units of Product AA. The total standard and actual costs for materials and direct labor for the 100 units of Product AA are as follows:

Materials: Standard Actual
Standard: 210 pounds at $3.00 per pound $630
Actual: 240 pounds at $2.85 per pound $684
Direct labor:
Standard: 400 hours at $15.00 per hour 6,000
Actual: 368 hours at $16.50 per hour 6,072


What is the material price variance for Bender Corporation?

a.$36 (F)

b.$36 (U)

c.$90 (F)

d.$30 (U)

Montana Company uses a standard costing system. The following information pertains to direct labor costs for the month of February:

Standard direct labor rate per hour $15.00
Actual direct labor rate per hour $13.50
Labor rate variance $18,000 favorable
Actual output 1,000 units
Standard hours allowed for actual production 10,000 hours


How many actual labor hours were worked during February for Montana Company?

a.2,000

b.10,000

c.1,200

d.12,000
Answer #1

1) Variable overhead efficiency variance = (2900*2-6000)*6 = 1200 U

So answer is b) $1200 U

2) Variable overhead price variance = (12*7000-90000) = 6000 U

So answer is a) $6000 U

3) Material price variance = (3000-2950)*18 = 900 F

So answer is a) $900 F

4) Material price variance = (3-2.85)*240 = 36 F

So answer is a) $36 F

5) Labor rate variance = (Standard rate-actual rate)actual hour

18000 = (15-13.50)X

Actual hour = 18000/1.5 = 12000

So answer is d) 12000